OpenStudy (anonymous):

If C = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12} and D = {numbers divisible by 3 from 1 to 12}, what is C ∩ D? A) {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12} B) {3, 6} C) {3, 6, 9, 12} D) { } I chose A, is that correct?

OpenStudy (anonymous):

@Notamathgenius @Notamathgenuis1 (not sure which one you are on .-.)

OpenStudy (notamathgenius):

Hai good morning, Ok You need to take all those numbers and see which ones divide EQUALLY into 3, I can say D is out. :D and so is A, now you do the rest .:D

OpenStudy (anonymous):

It is C. 4numbers between 1 and 12 are divisible by 3

OpenStudy (notamathgenius):

OMG IM teaching here Not giving Answers :(

OpenStudy (anonymous):

Ohhh sorry :)

OpenStudy (anonymous):

thanks @Notamathgenius for teaching me(: and thanks @Munqith xD

OpenStudy (notamathgenius):

haleyy don't look at his response >:D No problem :D d

OpenStudy (anonymous):

haha i shant! >;D

OpenStudy (anonymous):

Hhhh anytime

OpenStudy (notamathgenius):

Anymore?

OpenStudy (anonymous):

pftttttt..... of course!

OpenStudy (notamathgenius):

WOOHOO

OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5/3; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = -3/5; y-intercept = (0, 2) slope = 5/3; y-intercept = (0, -6)

OpenStudy (anonymous):

So guys I' new here, can anyone help me figure out how to use this site?

OpenStudy (notamathgenius):

Yea message me i will tell you. Haleyy Like yesterday, just plug in those numbers on the right for x and y and see which ones give you -6 :D

OpenStudy (anonymous):

ummm its pretty simple! lol you just click the people you wanna 'fan' 'block' 'view' or 'message' and then click your option! and to get someone to view what you want, just type '@' then their name. Ex: @Munqith

OpenStudy (anonymous):

Ahh okay soblet try messaging our NOT math genius :)

OpenStudy (anonymous):

oh but the not is silent! :D lol... you ARE a math genius c':

OpenStudy (notamathgenius):

haha thanks I feel special :P I am good at helping just not doing it myself :D

OpenStudy (anonymous):

haha im bad at all of it! :D lol help my last one please c:

OpenStudy (notamathgenius):

Ok I can tell you its either A or C not saying Which one ;D

OpenStudy (anonymous):

hmmmnmmm... imma guess..... A?!?!

OpenStudy (notamathgenius):

XDDDDD To funny, you sure?

OpenStudy (anonymous):

maybe e.e am i?

OpenStudy (notamathgenius):

^_^ hehe Just chose gurl.

OpenStudy (anonymous):

Hhh okay so this is gonna sound i'm on a mobile phone and it seems that the site isn't quite compatible with it, so everytime I try to click on "become a fan" the wind9w scrolls down because one of yu is typing

OpenStudy (anonymous):

....are you a boy or a girl? ?yoou give off girly vibes e.e @Notamathgenius

OpenStudy (notamathgenius):

I am a guy that likes guys lol

OpenStudy (anonymous):

oh cool! :D lol i like guys who like girls who are me xD

OpenStudy (notamathgenius):

I lean towards girls but i like guys :P

OpenStudy (anonymous):

you gotta boy fran? i gotta boyfran c:

OpenStudy (anonymous):

IS THE ANSWER C?!?

OpenStudy (notamathgenius):

No The guy i like has a bf :( But i have a gf :D

OpenStudy (notamathgenius):

OMG Finally Haleyy :P

OpenStudy (anonymous):

lol xD thats crazy! and yay!!!

OpenStudy (notamathgenius):

yes i must get him >:D

OpenStudy (anonymous):

get him hunny boo boo ;D lol i hate that fat ugly grimlin >.<

OpenStudy (anonymous):

not saying i hate you though o.o

OpenStudy (notamathgenius):

^_^ haha i will :P someday Hate me idc :P up to you

OpenStudy (anonymous):

nope nope never :p your helpful, i need you to be fran c:

OpenStudy (notamathgenius):

haha Ok :D be my friend then :D

OpenStudy (anonymous):

plus your cool as stuff n_n

OpenStudy (notamathgenius):

haha thank you, you are to funny though :D

OpenStudy (anonymous):

The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. d(x) = 1 + x/8 If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function? 0 ≤ x ≤ 8 0 ≤ x ≤ 24 0 ≤ x ≤ 40 1 ≤ x ≤ 8

OpenStudy (notamathgenius):

@TrashMan914 Need some assistance, mate.

OpenStudy (notamathgenius):

This is confusing, D:

OpenStudy (anonymous):

i know i forgot how to do it :(

OpenStudy (notamathgenius):

I never liked Domains, let me get Trash. hes a genius at this

OpenStudy (anonymous):

trash genius!! xD

OpenStudy (notamathgenius):

XDDD Like smarter than me :P

OpenStudy (anonymous):

and me! :D

OpenStudy (notamathgenius):

HAHA yes.

OpenStudy (anonymous):

i wanna say B e.e

OpenStudy (notamathgenius):

What about D?

OpenStudy (anonymous):

C ∩ D is whatever is common to both sets C and D. Set D contains all of the numbers divisible by 3 from 1 to 12 (which are 3, 6, 9, and 12). So: C = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12} and D = {3, 6, 9, 12}. Thus, C ∩ D = {3, 6, 9, 12}.

OpenStudy (notamathgenius):

We got that.

OpenStudy (anonymous):

lol. We are on a different question, it is the last one, hold on imma paste it here again

OpenStudy (anonymous):

The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. d(x) = 1 + x/8 If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function? 0 ≤ x ≤ 8 0 ≤ x ≤ 24 0 ≤ x ≤ 40 1 ≤ x ≤ 8

OpenStudy (anonymous):

I think it is B.

OpenStudy (notamathgenius):

Guessing? :P

OpenStudy (anonymous):

no? o.o

OpenStudy (notamathgenius):

haha ok ^_^

OpenStudy (anonymous):

depth varies from 1 ft to 4 ft 1 ≤ d ≤ 4 1 ≤ 1 + x/8 ≤ 4 .... multiply by 8 8 ≤ 8 + x ≤ 32 0 ≤ x ≤ 24

OpenStudy (anonymous):

1 ≤ d ≤ 4 1 ≤ 1 + x/8 ≤ 4 8 ≤ 8 + x ≤ 32 0 ≤ x ≤ 24

OpenStudy (notamathgenius):

OMG YOU WERE RIGH!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

what he said^^

OpenStudy (notamathgenius):

Right*

OpenStudy (anonymous):

be proud!

OpenStudy (notamathgenius):

I am very proud tear tear :"D

OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -5x + 15/-5 A) -x - 3 B) x + 15 C) x + 3 D) x - 3

OpenStudy (notamathgenius):

I know this!

OpenStudy (anonymous):

yay!!

OpenStudy (notamathgenius):

Well, I am thinking you may have to eliminate A.

OpenStudy (anonymous):

and B?

OpenStudy (anonymous):

...Well, for starters, 15/5 is 3...

OpenStudy (notamathgenius):

I think its D, Trash am I right?

OpenStudy (anonymous):

My calculations show -5x - 3, which is not an answer here.

OpenStudy (anonymous):

so how bout C? i think C.

OpenStudy (notamathgenius):

Well we know its going to have -5 so D is the most conclusive answer.

OpenStudy (anonymous):

@TrashMan914

OpenStudy (notamathgenius):

Haleyy, I would go with D

OpenStudy (anonymous):

D is the most logical.

OpenStudy (anonymous):

okay c:

OpenStudy (notamathgenius):

yes i was right >:D

OpenStudy (anonymous):

Layla has allotted $76 for gifts while on vacation. She plans to buy a souvenir for her parents for $16 and spend the rest on 4 T-shirts for her friends. If x represents the price, in dollars, of a T-shirt, which inequality correctly represents the money Layla will spend on souvenirs? 4x + 16 ≤ 76 16x + 4 ≤ 76 4x + 16 ≥ 76 16x + 4 ≥ 76

OpenStudy (anonymous):

A?

OpenStudy (notamathgenius):

Its going to have to be less that or equal to 76, so B Is wrong. and D is wrong so it leaves A or C ;D Chose wisely

OpenStudy (anonymous):

i think it is A though... is it right?! just tell meh c:

OpenStudy (notamathgenius):

No, Because it is saying its going to be Greater than or Equal to, and she only wants to spend 76 not more than that. SO its C

OpenStudy (anonymous):

oh .-.

OpenStudy (notamathgenius):

Don't be upset, mistakes happen :D